# Assessing the Uncertainty and Robustness of Object Detection Models for Detecting Stickers on Laptops

### v1

@@ 3.3 VLM-based Image Generation
VLMs are generic pre-trained multimodal models that learn from large-scale image-text pairs sourced from the Internet, enabling them to directly address downstream visual tasks without task-specific fine-tuning [21]. Therefore, we apply the zero-shot generalization ability of VLMs to generate a synthetic image dataset to study whether SDMs can identify stickers on laptops that they have not encountered before. Following the representative categories of VLMs, i.e., autoregressive and diffusion models [22], we select two state-of-the-art VLMs: DALL-E-3 [10] and Stable Diffusion-3 [11]. DALL-E-3 is the latest VLM developed by OpenAI, which leverages advanced transformer architectures to generate images from textual descriptions. Stable Diffusion-3, developed by StablilityAI, is a state-of-the-art text-to-image generation VLM that uses diffusion processes to create high-quality images from textual descriptions.
To ensure consistency and equity in using VLMs, we design the same prompt template to prompt the selected VLMs. Table 1 lists the main prompt phrases and relevant examples, which involve common laptop models, the number, size, style, and location of generated stickers, and other detailed image content information. The prompt template consists of these phrases, e.g., "a laptop, three stickers on the lid, one small music sticker, one small animal sticker, one small superhero sticker, stickers close to each other, top-down view, screen turn off, on a flat surface, a solid dark background". Note that to enhance the generality of the synthesized dataset, the prompt template omits information about the quality of the generated images, e.g., sharp focus or highly detailed, thereby generating images with varying levels of clarity that can be used to test and evaluate the generalization capability of SDMs.
@@ 4.1 Sticker Detection Models
We obtain six pre-trained SDMs from DTI, which represent state-of-the-art DNN-based object detection models. These models are designed by considering multiple model architectures which are Faster R-CNN [23], Faster R-CNN_v2 [24], RetinaNet [25], RetinaNet_v2 [26], SSD300 [27], and SSDLite [28]. DTI trained these models using a specially designed sticker detection dataset containing thousands of labeled laptop images with stickers in various poses, sizes, and lighting conditions. All six SDMs are trained based on the open-source implementation in PyTorch [29].
@@ 4.3 Evaluation Metrics
Mean Average Precision (mAP) is a commonly used metric for evaluating the accuracy of object detectors [30] and calculated as:
$\textit{mAP}=\frac{1}{nc}\sum_{c=1}^{nc}AP_{c},$ (13)
@@
Robustness Score (RS) measures the robustness of each SDM in performing the sticker detection task. We calculate RS regarding mAP and UQM, respectively, i.e., $RS_{mAP}$ and $RS_{uq}$ by considering the model’s performance on dataset $\mathcal{D}$ and $\mathcal{D}$’s adversarial version $\mathcal{D}_{adv}$. Specifically, for image $X\in\mathcal{D}$, there are 10 adversarial images $\{X_{adv}^{1},X_{adv}^{2},...,X_{adv}^{10}\}$ from $\mathcal{D}_{adv}$, and we consider RS regarding metric $v$ to be high when the model performs well in terms of $v$ (i.e., high mAP or low UQM) and the difference in $v$ between $X$ and $X$’s 10 adversarial examples is small. We then calculate $RS_{mAP}$ as:
$RS_{mAP}=Avg_{mAP}-Diff_{mAP},$ (14)
where $Avg_{mAP}$ is the mean value of mAP calculated by averaging the value achieved on $X$ and $X$’s 10 adversarial examples: $(mAP_{X}+\sum_{i=1}^{10}mAP_{X_{adv}^{i}})/11$; $Diff_{mAP}$ measures how differently the model performs on $X$ and $X$’s 10 adversarial examples in terms of mAP: $\sum_{i=1}^{10}|mAP_{X}-mAP_{X_{adv}^{i}}|/10$. As for $RS_{uq}$, we first calculate RS for each uncertainty metric $v$ as:
$RS_{v}=1-(Avg_{v}+Diff_{v}),$ (15)
where $v\in\{\textit{VR},\textit{SE},\textit{MI},\textit{TV},\textit{PS}\}$. $Avg_{v}$ and $Diff_{v}$ are calculated using the same equations as $Avg_{mAP}$ and $Diff_{mAP}$. We then calculate $RS_{uq}$ by taking the mean value of RS for each uncertainty metric:
$RS_{uq}=\frac{\sum_{v\in UQMs}RS_{v}}{count(UQMs)}.$ (16)
@@ 4.5 Statistical Test
Since we compare more than two paired groups, i.e., SDMs using the same image dataset, we conduct the statistical analysis using the Friedman test [31], the Wilcoxon Signed-Rank test [32], and rank-biserial correlation [33]. Following the guideline [34, 35], we employ the Friedman test to verify if there are overall significant differences among all the paired groups (i.e., the models). If at least one group differs significantly from the others, $p<0.05$ will be computed by this statistical test. If this is the case, we apply the Wilcoxon Signed-Rank test for pairwise comparisons (i.e., comparing two models) with a significance level of 5%. To interpret the Wilcoxon Signed-Rank test more precisely, we calculate the median and the standard deviation by the median absolute deviation (MAD) of our data, i.e., [median $\pm$ 2 MAD] [36, 37]. It helps to check if the underlying distribution is stochastically equal to, less than, or greater than a distribution symmetric about zero, and accordingly define the relevant alternative hypothesis for more effective execution of the Wilcoxon Signed-Rank test. Moreover, for pairwise groups with significant differences, we use the rank-biserial correlation as the effect size of the Wilcoxon Signed-Rank test to determine which group shows better results. The rank-biserial correlation ranges from -1 to 1, where a positive result indicates that the first group tends to be larger than the second, and a negative result means the opposite, whereas 0 explains there is no significant difference between the two paired groups. Based on the guideline [38, 39], the effect size magnitude of the rank-biserial correlation involves four levels: negligible ($>$ 0.0 and $<$ 0.1), small ($\geq$ 0.1 and $<$ 0.3), medium ($\geq$ 0.3 and $<$ 0.5), and large ($\geq$ 0.5 and $\leq$ 1.0). We consider medium and large levels to demonstrate significant differences in our data.
To study the correlation between mAP and UQMs for RQ4, we use Spearman’s rank correlation coefficient [40], which provides the correlation coefficient ($\rho$) and the significance level ($p$). A $p$ less than 0.05 indicates a significant correlation between mAP and UQMs. Whereas correlation coefficient $\rho$ ranges from -1 to 1, revealing the direction and strength of the correlation. A positive $\rho$ indicates a positive correlation, meaning that one variable increases as the other increases, and vice versa. A $\rho$ of 0 shows no correlation.
Considering that the Wilcoxon Signed-Rank test and Spearman’s rank correlation coefficient are utilized multiple times to compare multiple paired groups, we address the issue of aggregated error probabilities resulting from multiple comparisons by applying the Holm–Bonferroni method [41]. This multiple comparison correction technique adjusts the significance level from the overall and controls the family-wise error rate at an $\alpha$ level, e.g., 5%.
@@ 5 Results and Analyses
We present descriptive statistics and summarize the key results of the statistical tests in Sections 5.1 to 5.4, and then provide guidelines for selecting SDMs in Section 5.5. The complete statistical results for each RQ are available in our replication package [42].
@@ 5.2 Results for RQ2 – Uncertainty
Figure 4 presents the mean results regarding the five UQMs for each dataset at different dropout rates with 95% confidence intervals.
@@
Uncertainty for Bounding Box Regression. Similar observations can be made for the two bounding box regressions UQMs (i.e., TV and PS), that is, SSD300 and SSDLite show higher TV and PS values for all datasets at all dropout rates except for TV on dalleImg dataset, where SSD300 achieves lower TV than Faster R-CNN until the dropout rate reaches 0.45. Besides, Faster R-CNN, Faster R-CNN_v2, RetinaNet, and RetinaNet_v2 achieve comparable and consistent low TV and PS, except for TV of Faster R-CNN on dalleImg dataset. The results of TV and PS show that SSD300 and SSDLite exhibit higher uncertainty in predicting bounding boxes than the other four SDMs. Moreover, when the dropout rate increases, the uncertainties in SSD300 and SSDLite also increase, while for the other SDMs, the dropout rate has little effect on their uncertainties.
Table 3 shows the best SDM(s), where we can find that regarding VR, Faster R-CNN, Faster R-CNN_v2, SSD300, and SSDLite achieve the best performance on origImg dataset, while all SDMs perform comparable on dalleImg and stableImg datasets. As for SE, MI, TV, and PS, RetinaNet_v2 perform the best, i.e., shows the lowest uncertainty when detecting stickers, except for SE on origImg dataset, where Faster R-CNN and RetinaNet_v2 achieve comparable best performance. When looking at the SDM recommendations across all datasets, regarding VR, Faster R-CNN, Faster R-CNN_v2, SSD300, and SSDLite are recommended, while for SE, MI, TV, and PS, RetinaNet_v2 is recommended as the best SDM.
@@
Figure 5: Average $\bm{\mathit{RS_{mAP}}}$ and $\bm{\mathit{RS_{uq}}}$ achieved by SDMs for origImg, dalleImg, and stableImg at different dropout rates. M1: Faster R-CNN, M2: Faster R-CNN_v2, M3: RetinaNet, M4: RetinaNet_v2, M5: SSD300, and M6: SSDLiteRQ3.
@@ 5.4 Results for RQ4 – Correlation
Table 4: Results of the Spearman’s rank correlation coefficient between mAP and UQMs for each SDM on all dropout rates and datasets. A value or N/A denotes that the correlation is not statistically significant or non-existent, otherwise, it is significant, i.e., $\bm{p<0.05}$RQ4.
@@ 5.5 Concluding Remarks and Guidelines
We analyze the performance of each SDM in terms of detection accuracy (mAP), prediction uncertainty (UQM), and adversarial robustness ($RS_{mAP}$ and $RS_{uq}$). Based on the results, we observe that different SDMs perform differently regarding different metrics. In practice, considering different purposes and applications, selecting the appropriate SDM based on specific performance requirements is crucial. Therefore, we provide guidelines for selecting SDMs from various perspectives in Figure 6. Regarding detection accuracy, i.e., mAP, Faster R-CNN_v2 is recommended as it achieves the overall best performance. As for UQMs, for both label classification and bounding box regression tasks, we recommend RetinaNet_v2 because it shows lower uncertainty in terms of both types of uncertainty metrics. When it comes to the robustness of SDMs, Faster R-CNN_v2 shows the overall highest robustness in terms of detection accuracy, i.e., $RS_{mAP}$, and therefore recommended as the best SDM, while when considering robustness regarding uncertainty, RetinaNet_v2 is recommended as the best SDM.
@@ 6 Threats to Validity
Conclusion Validity concerns the reliability of the conclusions. We employed appropriate statistical tests to draw reliable conclusions and followed a rigorous statistical procedure to analyze the collected data. Internal Validility is related to the parameter settings. To determine the hyperparameters for the clustering algorithm, i.e., HDBSCAN, we ran different combinations of hyperparameter values on a small subset of the original dataset and selected the combination that correctly clustered all objects. Besides, the threshold value $t$ for IoU is set to 0.5, following Catak et al. [20]. We, however, acknowledge that experimenting with various $t$ is needed to study its effect on the results, which is one of our further plans. Constructive Validity concerns the metrics used for the evaluation and we employed comparable metrics to ensure fair comparisons. Specifically, we used mAP to measure the sticker detection performance and five UQMs commonly used in classification and regression to measure model uncertainty. We also calculated RS to measure the robustness of the models under adversarial attacks. External Validility is about the generalizability of the empirical evaluation. We employ six models and six datasets. These models are based on different architectural designs and represent state-of-the-art object detection models. Regarding the benchmark datasets, in addition to the original dataset, we synthesize two datasets by prompting two state-of-the-art VLMs, i.e., DALL-E-3 and Stable Diffusion-3. Besides, we also generate adversarial datasets by perturbing the original and synthetic datasets.
@@ 7 Lessons Learned and Industrial Perspectives
Understanding, Utilizing, and Handling Uncertainty in Refurbishing Laptops. We quantify uncertainty in the SDMs from two aspects: uncertainty in label classification and bounding box regression, where the former tells the confidence of SDMs in the classification of detected stickers, whereas the latter represents their confidence and variability in localizing the positions of the detected stickers. UQ plays a pivotal role in enhancing the refurbishment of laptops by guiding laptop image data collection, ensuring dataset diversity, and determining the need for manual intervention. Collecting laptop images is time-consuming and significantly impacts solution quality, and UQ can guide decisions on whether additional data is needed. Specifically, UQ helps identify areas with high prediction uncertainty, indicating the need for more diverse and comprehensive training data. Besides, UQ is essential in addressing the limitations of existing datasets created by DTI, which primarily features Lenovo laptops and common sticker placements, e.g., stickers on the back cover. Also, images in the DTI dataset were captured in a relatively sterile setup with high quality, which may not reflect real-world conditions. Therefore, SDMs need to be tested under different conditions in terms of handling unexpected situations to ensure robustness, such as stickers on the side of a laptop or stickers near a mouse pad, where processor branding stickers are often found. By highlighting gaps in data diversity, UQ directs efforts towards collecting images from various laptop brands and unexpected sticker locations, ensuring SDMs are robust and adaptable to real-world conditions. Finally, UQ informs the necessity of manual intervention in the refurbishment process, enabling efficient and reliable operations by indicating when human oversight is required.
In addition to the uncertainty in SDMs, the overall process of laptop refurbishing experiences uncertainties from various aspects, such as humans working in collaboration with robots, software deployed in robots, hardware errors, and other environmental factors (e.g., lighting conditions). Hence, various UQ methods are needed to capture uncertainties in the whole process, not only related to the employed SDMs. As a result, an integrated UQ method is required for the entire laptop refurbishing process that quantifies uncertainties in various aspects and eventually provides an overall uncertainty associated with the whole process. Such quantified uncertainty can play a key role in identifying factors contributing to uncertainties, followed by devising best practices and methods to reduce the overall uncertainty.
@@
Realism of VLM-Generated Images. To assess the SDMs, we generated images of laptops with stickers using two commonly used pre-trained VLMs. However, this poses an open question of whether the images generated are realistic. In our context, all the generated images were manually checked for realism. However, such an approach is not scalable, especially when generating a large-scale dataset. To this end, we foresee the need for an automated method to generate realistic images of laptops with stickers. Pre-trained VLMs already possess a strong understanding of general image features, and domain-specific fine-tuning can further refine their output. Therefore, one possible solution is to fine-tune VLMs to adapt to specific domains related to laptops and stickers, thereby helping produce more realistic images from the outset and reducing the need for extensive post-production filtering. Besides, large language models have been proven to be a way of realism evaluation [43]. Furthermore, training a realism classifier to distinguish realistic from unrealistic images can automate the manual check process. In the future, we plan to investigate scalable and efficient methods to generate large-scale realistic laptop image datasets.

### v2

@@ 3.3 VLM-based Image Generation
VLMs are generic pre-trained multimodal models that learn from large-scale image-text pairs sourced from the Internet, enabling them to directly address downstream visual tasks without task-specific fine-tuning [17]. Therefore, we apply the zero-shot generalization ability of VLMs to generate a synthetic image dataset to study whether SDMs can identify stickers on laptops that they have not encountered before. Following the representative categories of VLMs, i.e., autoregressive and diffusion models [18], we select two state-of-the-art VLMs: DALL-E-3 [5] and Stable Diffusion-3 [6]. DALL-E-3 is the latest VLM developed by OpenAI, leveraging advanced transformer architectures to generate images from textual descriptions. Stable Diffusion-3, developed by StablilityAI, is a text-to-image generation VLM that uses diffusion processes to create images from textual descriptions.
To ensure consistency and equity in using VLMs, we design the same prompt template to prompt the selected VLMs. Table 1 lists the main prompt phrases and relevant examples, which involve common laptop models, the number, size, style, and location of generated stickers, and other detailed image content information. The prompt template consists of these phrases, e.g., “a laptop, three stickers on the lid, one small music sticker, one small animal sticker, one small superhero sticker, stickers close to each other, top-down view, screen turn off, on a flat surface, a solid dark background". Note that to enhance the generality of the synthesized dataset, the prompt template omits information about the quality of the generated images, e.g., sharp focus or highly detailed, thereby generating images with varying levels of clarity that can be used to test and evaluate the generalization capability of SDMs.
@@ 4.1 Sticker Detection Models
We obtain six pre-trained SDMs employed in the robotic software, representing state-of-the-art DNN-based object detection models. These models are designed by considering multiple model architectures which are Faster R-CNN [20], Faster R-CNN_v2 [21], RetinaNet [22], RetinaNet_v2 [23], SSD300 [24], and SSDLite [25]. DTI trained these models using a specially designed sticker detection dataset containing thousands of labeled laptop images with stickers in various poses, sizes, and lighting conditions. All six SDMs are trained based on the open-source implementation in PyTorch [26].
@@ 4.3 Evaluation Metrics
Mean Average Precision (mAP) is a metric for evaluating the accuracy of object detectors [27] and is calculated as:
$\vspace{-2pt}\textit{mAP}=\frac{1}{nc}\sum_{c=1}^{nc}AP_{c},$ (13)
@@
Robustness Score (RS) measures the robustness of each SDM in performing the sticker detection task. We calculate RS regarding mAP and UQM, respectively, i.e., $RS_{mAP}$ and $RS_{uq}$ by considering the model’s performance on dataset $\mathcal{D}$ and $\mathcal{D}$’s adversarial version $\mathcal{D}_{adv}$. Specifically, for image $X\in\mathcal{D}$, there are 10 adversarial images $\{X_{adv}^{1},X_{adv}^{2},...,X_{adv}^{10}\}$ from $\mathcal{D}_{adv}$, and we consider RS regarding metric $\mathcal{M}$ to be high when the model performs well in terms of $\mathcal{M}$ (i.e., high mAP or low UQM) and the difference in $\mathcal{M}$ between $X$ and $X$’s 10 adversarial examples is small. We then calculate $RS_{mAP}$ as:
$\vspace{-2pt}RS_{mAP}=Avg_{mAP}-Diff_{mAP},$ (14)
where $Avg_{mAP}$ is the mean value of mAP calculated by averaging the value achieved on $X$ and $X$’s 10 adversarial examples: $(mAP_{X}+\sum_{i=1}^{10}mAP_{X_{adv}^{i}})/11$; $Diff_{mAP}$ measures how differently the model performs on $X$ and $X$’s 10 adversarial examples in terms of mAP: $\sum_{i=1}^{10}|mAP_{X}-mAP_{X_{adv}^{i}}|/10$. For $RS_{uq}$, we first calculate RS for each uncertainty metric $\mathcal{M}$:
$\vspace{-2pt}RS_{\mathcal{M}}=1-(Avg_{\mathcal{M}}+Diff_{\mathcal{M}}),$ (15)
where $\mathcal{M}\in\{\textit{VR},\textit{SE},\textit{MI},\textit{TV},\textit{PS}\}$. $Avg_{\mathcal{M}}$ and $Diff_{\mathcal{M}}$ are calculated using the same equations as $Avg_{mAP}$ and $Diff_{mAP}$. We then calculate $RS_{uq}$ by taking the mean value of RS for each uncertainty metric:
$\vspace{-1pt}RS_{uq}=\frac{\sum_{\mathcal{M}\in UQMs}RS_{\mathcal{M}}}{count(UQMs)}.$ (16)
@@ 4.5 Statistical Test
Since we compare more than two paired groups, i.e., SDMs using the same image dataset, we conduct the statistical analysis using the Friedman test [28], the Wilcoxon Signed-Rank test [29], and rank-biserial correlation [30]. Following the guideline [31, 32], we employ the Friedman test to verify if there are overall significant differences among all the paired groups (i.e., the models). If at least one group differs significantly from the others, $p<0.05$ will be computed by this statistical test. If this is the case, we apply the Wilcoxon Signed-Rank test for pairwise comparisons (i.e., comparing two models) with a significance level of 5%. To interpret the Wilcoxon Signed-Rank test more precisely, we calculate the median and the standard deviation by the median absolute deviation (MAD) of our data, i.e., [median $\pm$ 2 MAD] [33, 34]. It checks if the underlying distribution is stochastically equal to, less than, or greater than a distribution symmetric about zero, and accordingly defines the relevant alternative hypothesis for more effective execution of the Wilcoxon Signed-Rank test. Moreover, for pairwise groups with significant differences, we use the rank-biserial correlation as the effect size of the Wilcoxon Signed-Rank test to determine which group shows better results. The rank-biserial correlation ranges from -1 to 1, where a positive result indicates that the first group tends to be larger than the second, and a negative result means the opposite, whereas 0 explains there is no significant difference between the two paired groups.
To study the correlation between mAP and UQMs for RQ4, we use Spearman’s rank correlation coefficient [35], which provides the correlation coefficient ($\rho$) and the significance level ($p$). A $p$ less than 0.05 indicates a significant correlation between mAP and UQMs, whereas $\rho$ ranges from -1 to 1, revealing the direction and strength of the correlation. A positive $\rho$ indicates a positive correlation, meaning that one variable increases as the other increases, and vice versa. A $\rho$ of 0 shows no correlation. Considering that the Wilcoxon Signed-Rank test and Spearman’s rank correlation coefficient are utilized multiple times to compare multiple paired groups, we address the issue of aggregated error probabilities resulting from multiple comparisons by applying the Holm–Bonferroni method [36]. This multiple comparison correction technique adjusts the significance level from the overall and controls the family-wise error rate at an $\alpha$ level, e.g., 5%.
@@ 5 Results and Analyses
We present descriptive statistics and summarize the key results of the statistical tests in Sections 5.1 to 5.4, and then provide guidelines for selecting SDMs in Section 5.5. The complete results are available in our replication package [37].
@@ 5.2 Results for RQ2 – Uncertainty
Figure 4 presents the mean results of UQMs for each dataset at different dropout rates with 95% confidence intervals.
@@
Uncertainty for Bounding Box Regression. Similar observations can be made for the bounding box regression UQMs, that is, SSD300 and SSDLite show higher TV and PS values for all datasets at all dropout rates except for TV on dalleImg dataset, where SSD300 achieves lower TV than Faster R-CNN until the dropout rate reaches 0.45. Besides, Faster R-CNN, Faster R-CNN_v2, RetinaNet, and RetinaNet_v2 achieve comparable and consistent low TV and PS, except for TV of Faster R-CNN on dalleImg dataset. The results of TV and PS show that SSD300 and SSDLite exhibit higher uncertainty in predicting bounding boxes than the other four SDMs. Moreover, when the dropout rate increases, the uncertainties in SSD300 and SSDLite also increase, while for the other SDMs, the dropout rate has little effect on their uncertainties.
Table 3 shows the best SDM(s), where we can find that regarding VR, Faster R-CNN, Faster R-CNN_v2, SSD300, and SSDLite achieve the best performance on origImg dataset, while all SDMs perform comparable on dalleImg and stableImg datasets. For SE, MI, TV, and PS, RetinaNet_v2 perform the best, i.e., shows the lowest uncertainty, except for SE on origImg dataset, where Faster R-CNN and RetinaNet_v2 achieve comparable best performance. When looking at the SDM recommendations across all datasets, regarding VR, Faster R-CNN, Faster R-CNN_v2, SSD300, and SSDLite are recommended, while for SE, MI, TV, and PS, RetinaNet_v2 is recommended as the best SDM.
@@
Figure 5: Average $RS_{mAP}$ and $RS_{uq}$ achieved by SDMs for origImg, dalleImg, and stableImg at different dropout rates. M1: Faster R-CNN, M2: Faster R-CNN_v2, M3: RetinaNet, M4: RetinaNet_v2, M5: SSD300, and M6: SSDLiteRQ3.
@@ 5.4 Results for RQ4 – Correlation
Table 4: Results of the Spearman’s rank correlation coefficient between mAP and UQMs for each SDM on all dropout rates and datasets. A value or N/A denotes that the correlation is not statistically significant or non-existent, otherwise, it is significant, i.e., $p<0.05$RQ4.
@@ 5.5 Concluding Remarks and Guidelines
We analyze the performance of each SDM regarding detection accuracy (mAP), prediction uncertainty (UQM), and adversarial robustness ($RS_{mAP}$ and $RS_{uq}$). Based on the results, we observe that different SDMs perform differently regarding different metrics. In practice, considering different purposes and applications, selecting the appropriate SDM based on specific performance requirements is crucial. Therefore, we provide guidelines for selecting SDMs from various perspectives in Figure 6. Regarding detection accuracy, i.e., mAP, Faster R-CNN_v2 is recommended as it achieves the overall best performance. As for UQMs, for both label classification and bounding box regression tasks, we recommend RetinaNet_v2 as it shows lower uncertainty in terms of both types of uncertainty metrics. For the robustness of SDMs, Faster R-CNN_v2 shows the overall highest robustness regarding detection accuracy, i.e., $RS_{mAP}$, and is recommended as the best SDM. When considering robustness regarding uncertainty, RetinaNet_v2 is recommended as the best SDM. Based on the guidelines, adaptive software integrated within the robot to select SDMs would significantly enhance the robot’s decision-making capabilities, particularly in dynamic and uncertain environments. Such software can automatically select the most appropriate SDMs by considering various metrics.
@@ 6 Threats to Validity
Conclusion Validity concerns the reliability of the conclusions. We employed appropriate statistical tests to draw reliable conclusions and followed a rigorous statistical procedure to analyze the collected data. Internal Validity is related to the parameter settings. To determine the hyperparameters for the clustering algorithm, i.e., HDBSCAN, we ran different combinations of hyperparameter values on a small subset of the original dataset and selected the combination that correctly clustered all objects. Besides, the threshold value for IoU is set to 0.5, following Catak et al. [16]. Constructive Validity concerns the metrics used for the evaluation and we employed comparable metrics to ensure fair comparisons. Specifically, we used mAP to measure the sticker detection performance and five UQMs commonly used in classification and regression to measure model uncertainty. We also calculated RS to measure the robustness of the models under adversarial attacks. External Validility is about the generalizability of the empirical evaluation. We employ six models and six datasets. These models are based on different architectural designs and represent state-of-the-art object detection models. Regarding the benchmark datasets, in addition to the original dataset, we synthesize two datasets by prompting two VLMs, i.e., DALL-E-3 and Stable Diffusion-3. Besides, we generate adversarial datasets by perturbing the original and synthetic datasets.
@@ 7 Lessons Learned and Industrial Perspectives
Understanding, Utilizing, and Handling Uncertainty in Refurbishing Laptop software. UQ plays a pivotal role in enhancing the refurbishment of laptops by guiding laptop image data collection, ensuring dataset diversity, and determining the need for manual intervention. Collecting laptop images is time-consuming and significantly impacts solution quality, and UQ can guide decisions on whether additional data is needed. Specifically, UQ helps identify areas with high prediction uncertainty, indicating the need for more diverse and comprehensive training data. Besides, UQ is essential in addressing the limitations of existing datasets created by DTI, which primarily features Lenovo laptops and common sticker placements, e.g., stickers on the back cover. Also, images in the DTI dataset were captured in a relatively sterile setup with high quality that may not reflect real-world conditions. Thus, SDMs need to be tested under different conditions in handling unexpected situations, such as stickers near a mouse pad, where processor branding stickers are often found. By highlighting gaps in data diversity, UQ directs efforts towards collecting images from various laptop brands and unexpected sticker locations, ensuring SDMs are robust and adaptable to real-world conditions. Finally, UQ informs the need for manual intervention in the refurbishment process, enabling efficient and reliable operations by indicating when human oversight is required. In addition to the uncertainty in SDMs, the overall process of laptop refurbishing experiences uncertainties from various aspects, such as humans working in collaboration with robots, hardware errors, and other environmental factors. Hence, a holistic UQ method is required for the entire laptop refurbishing process that quantifies uncertainties from various aspects and eventually provides an overall uncertainty. Such quantified uncertainty plays a key role in identifying factors contributing to uncertainties, followed by devising best practices to reduce the overall uncertainty.
@@
Realism of VLM-Generated Images. To assess the SDMs, we generated images of laptops with stickers using two commonly used pre-trained VLMs. However, this poses an open question of whether the images generated are realistic. In our context, all the generated images were manually checked for realism. However, such an approach is not scalable, especially when generating a large-scale dataset. To this end, we foresee the need for an automated method to generate realistic images of laptops with stickers. Pre-trained VLMs already possess a strong understanding of general image features, and domain-specific fine-tuning can further refine their output. Thus, one possible solution is to fine-tune VLMs to adapt to specific domains related to laptops and stickers, thereby helping produce more realistic images from the outset and reducing the need for extensive post-production filtering. Besides, large language models can be another way of realism evaluation [38]. Furthermore, training a realism classifier to distinguish realistic from unrealistic images can automate the manual check process. In the future, we plan to investigate scalable and efficient methods to generate large-scale realistic laptop image datasets.